OpenStudy (anonymous):

@jim_thompson5910 4. Graph the circle with center at (0, 4) and radius 2. Write the equation of the circle.

jimthompson5910 (jim_thompson5910):

In this case h = 0 k = 4 r = 2

OpenStudy (anonymous):

can you do the equation for me please :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

start with (x-h)^2 + (y-k)^2 = r^2 and plug in h = 0, k = 4 and r = 2

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

(x-0)^2+(y-5)^2=2^2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

k = 4, not k = 5

OpenStudy (anonymous):

oh *y-4

jimthompson5910 (jim_thompson5910):

good, and 2^2 = 4 finally x-0 is the same as just x

jimthompson5910 (jim_thompson5910):

so (x-0)^2+(y-4)^2=2^2 turns into x^2+(y-4)^2=4